# Compositional–ARC: Assessing Systematic Generalization in Abstract Spatial Reasoning

### v2

@@ A.4 Dataset statistics
Appendix B Training details
@@
Training is conducted over 200 epochs with a batch size of 200 for the standard few-shot learning task (i.e., $200\cdot 10=2000$ queries per batch), and over 300 epochs with the same batch size for the systematicity task. A learning rate of 0.01 is used in both cases. Following the approach of Lake & Baroni (2023), we apply a warm-up phase during the first episode, beginning with a learning rate of $1\times 10^{-4}$, followed by a linear decay to $5\times 10^{-4}$ over the course of training. Additional hyperparameter settings are provided in Section B.1 and summarized in Table 3.
@@ B.2 Implementation details
Appendix C Experiment details
This section provide further details regarding our experimental setup. Specifically, Section C.1 presents formal definitions of the evaluation metrics used to assess the performance of the models studied in this work, while Section C.2 outlines additional information on how we interact with API-based LLMs.
C.1 Evaluation metrics
As described in Section 4.3, we use three different evaluation metrics to assess model performance in this study: i) exact match accuracy, ii) color accuracy, and iii) shape accuracy. These metrics are formally defined based on the grid-based environment $\bm{X}$} and the concept of an object $\mathbb{O}$}, as specified in Definition A.1.
@@ C.2 Model information
As described in Section 4.2, we evaluate three different general-purpose LLMs on Compositional-ARC. Specifically, we assess the performance of o3-mini (OpenAI, 2025) (version o3-mini-2025-01-31222https://platform.openai.com/docs/models/o3-mini), GPT-4o (Achiam et al., 2023) (version gpt-4o-2024-08-06333https://platform.openai.com/docs/models/gpt-4o), and Gemini 2.0 Flash (DeepMind, 2024) (version gemini-2.0-flash-001444https://ai.google.dev/gemini-api/docs/models#gemini-2.0-flash). All models are accessed via their respective batch APIs, enabling us to process multiple samples per request. Unless otherwise specified, we employ the default API settings. For GPT-4o and o3-mini, this corresponds to a temperature and top_p value of 1.01.0.555https://platform.openai.com/docs/api-reference/chat/create Due to financial constraints, the o3-mini model is configured with a “low” reasoning effort. For Gemini 2.0 Flash, the provider does not disclose default parameter settings.
@@
Prompts.
The complete set of prompts used in our evaluations is presented in Figures 11 through 14. To ensure consistency and facilitate meaningful comparisons, we apply the same prompts across all models. The standard few-shot learning prompt appears in Figure 11, while the prompt used for the systematicity task is shown in Figure 13. For Gemini 2.0 Flash, we add the instruction: “Do not generate any code to solve the task” to the output requirements, as the model otherwise does not adhere to the required output format. As outlined in Section 4.2, we additionally evaluate GPT-4o and Gemini 2.0 Flash in a multimodal configuration, in which both an image of the study examples and the input query are provided alongside the text prompt (text+image). The multimodal prompt for the few-shot learning task is shown in Figure 12, with the accompanying image illustrated in Figure 9. The corresponding multimodal prompt for the systematicity task is depicted in Figure 14, with the associated image presented in Figure 10. For the textual prompts, we represent grids as two-dimensional arrays, consistent with prior work (Moskvichev et al., 2023)). For instance, the final query input grid in Figure 4 would be represented as:
[…]
Domain-specific LLMs.
Appendix D Additional results
In this section, we present additional results for the experiments conducted in this study. First, we present additional qualitative results related to the model predictions on the standard few-shot learning and the systematicity task. Figures 4 through 6 illustrate representative episodes from the standard few-shot learning task. Model predictions are shown adjacent to each query, with results for GPT-4o and Gemini 2.0 Flash corresponding to text-only prompts. Across all three episodes, the model trained using MLC consistently predicts the correct output grid. In contrast, GPT-4o and Gemini 2.0 Flash frequently fail to identify the correct transformation—either misrepresenting the shape of the transformed object or incorrectly predicting its final position. Notably, o3-mini successfully predicts the correct output for the episodes in Figures 5 and 6, but fails on the example in Figure 4. Figures 7 and 8 highlight episodes from the systematicity task. As shown, all general-purpose LLMs fail to produce accurate transformations, often misplacing the transformed object within the grid. In contrast, the model trained via MLC consistently predicts the correct transformation.
@@ Appendix D Additional results
As outlined in Section C.2, the general-purpose LLMs we evaluate are instructed to present their final output grid predictions using the keyword “output:”, followed by a two-dimensional array of size $10\times 10$ in the format $\ldots$]]”. Responses that do not conform to this expected pattern are excluded from subsequent analyses and are not included in accuracy calculations. Table 4 provides an overview of the proportion of valid responses for each model. In the standard few-shot learning setting, all models demonstrate very high valid response rates, exceeding 99%. However, in the systematicity task, a slight decrease in valid responses is observed for Gemini 2.0 Flash when additional visual input (text+image) is introduced, with the rate falling to 94.09%. More significantly, GPT-4o exhibits a notable drop in valid response rate to 77.24% under multimodal conditions. We hypothesize that this decline may be attributed to the increased context length resulting from the additional image input.
@@
In addition to the model trained via MLC on a stream of dynamically changing visual interpretation grammars, as described in Section 3.2, we adopt the approach of Lake (2019) and train a transformer-based encoder-decoder on a dataset governed by a fixed visual grammar (referred to as basic seq2seq). This means that the indicator-transformation mappings are static across the whole dataset. For instance, if yellow object translates one step downward, then this applies to all data samples across the dataset. Instead of episodes with few-shot examples, this dataset comprises individual input-output grid pairs, where the objective is to predict the output grid corresponding to a given input grid. This more closely resembles a standard training approach.
@@
The model is trained for 200 epochs on the dataset using the parameters specified in Section B. While it successfully fits the training data (with an accuracy of over 99%), it fails to generalize to the out-of-distribution test set, achieving a test accuracy of 0.0%. This demonstrates that traditional model training, sample by sample, does not encourage systematic generalization to unseen compositions. Instead, systematicity requires a training procedure with examples over dynamically varying interpretation grammars, as described in Section 3.2.
Appendix E Use of AI assistants

### v3

@@ B.4 Dataset statistics
In a similar vein, Table 8 shows the statistics for the dataset version that includes more diverse transformations, as described in Section 5.3. The table provides information on the query transformation compositions present in the training and test sets, along with the frequency of each geometric transformation within the training dataset.
Appendix C Training details
@@
Training is conducted over 200 epochs with a batch size of 200 for the standard few-shot learning task (i.e., $200\cdot 10=2000$ queries per batch), and over 300 epochs with the same batch size for the systematicity task. A learning rate of 0.01 is used in both cases. Following the approach of Lake and Baroni (2023), we apply a warm-up phase during the first episode, beginning with a learning rate of $1\times 10^{-4}$, followed by a linear decay to $5\times 10^{-4}$ over the course of training. Additional hyperparameter settings are provided in Section C.1 and summarized in Table 3.
@@ C.2 Implementation details
C.3 Original MLC training
In the original MLC setup, Lake and Baroni (2023) train a standard seq2seq transformer (3-layer encoder/decoder, 8 attention heads, hidden dimension 128) with Adam on 100,000 dynamically generated episodes of pseudo-language instructions. Each episode is defined by a latent compositional grammar and contains multiple study examples and queries concatenated into a single input sequence, as described in Section 2. Training minimizes token-level cross-entropy for 50 epochs using a batch size of 25 episodes, a learning rate of $10^{-3}$ with a one-epoch warm-up followed by linear decay, and dropout of 0.1. Compared to our approach, their model targets linguistic sequences rather than visual grids and does not use patch-wise losses or background reweighting. Additionally, we train longer with larger batches and a loss designed to emphasize non-background spatial structure.
Appendix D Experiment details
This section provides further details regarding our experimental setup. Specifically, Section D.1 presents formal definitions of the evaluation metrics, while Section D.2 outlines additional information on how we interact with API-based LLMs.
D.1 Evaluation metrics
As described in Section 4.3, we use three different evaluation metrics to assess model performance: i) exact match accuracy, ii) color accuracy, and iii) shape accuracy. These metrics are formally defined based on the grid-based environment $\bm{X}$} and the concept of an object $\mathbb{O}$}, as specified in Definition B.1.
@@ D.2 Model information
As described in Section 4.2, we evaluate three different general-purpose LLMs on Compositional-ARC. Specifically, we assess the performance of o3-mini (OpenAI, 2025) (version o3-mini-2025-01-31333https://platform.openai.com/docs/models/o3-mini), GPT-4o (Achiam et al., 2023) (version gpt-4o-2024-08-06444https://platform.openai.com/docs/models/gpt-4o), and Gemini 2.0 Flash (DeepMind, 2024) (version gemini-2.0-flash-001555https://ai.google.dev/gemini-api/docs/models#gemini-2.0-flash). All models are accessed via their respective batch APIs, enabling us to process multiple samples per request. Unless otherwise specified, we employ the default API settings. For GPT-4o and o3-mini, this corresponds to a temperature and top_p value of 1.01.0.666https://platform.openai.com/docs/api-reference/chat/create Due to financial constraints, the o3-mini model is configured with a “low” reasoning effort. For Gemini 2.0 Flash, the provider does not disclose default parameter settings.
Prompts.
The complete set of prompts used in our evaluations is presented in Figures 12 through 15. To ensure consistency and facilitate meaningful comparisons, we apply the same prompts across all models. The standard few-shot learning prompt appears in Figure 12, while the prompt used for the systematicity task is shown in Figure 14. For Gemini 2.0 Flash, we add the instruction: “Do not generate any code to solve the task” to the output requirements, as the model otherwise does not adhere to the required output format. As outlined in Section 4.2, we additionally evaluate GPT-4o and Gemini 2.0 Flash in a multimodal configuration, in which both an image of the study examples and the input query are provided alongside the text prompt (text+image). The multimodal prompt for the few-shot learning task is shown in Figure 13, with the accompanying image illustrated in Figure 10. The corresponding multimodal prompt for the systematicity task is depicted in Figure 15, with the associated image presented in Figure 11. For the textual prompts, we represent grids as two-dimensional arrays, consistent with prior work (Moskvichev et al., 2023)). For instance, the final query input grid in Figure 5 would be represented as:
[…]
Domain-specific LLMs.
Appendix E Additional results
In this section, we present additional results for the experiments conducted in this study. First, we present additional qualitative results related to the model predictions on the standard few-shot learning and the systematicity task. Figures 5 through 7 illustrate representative episodes from the standard few-shot learning task. Model predictions are shown adjacent to each query, with results for GPT-4o and Gemini 2.0 Flash corresponding to text-only prompts. Across all three episodes, the model trained using MLC consistently predicts the correct output grid. In contrast, GPT-4o and Gemini 2.0 Flash frequently fail to identify the correct transformation—either misrepresenting the shape of the transformed object or incorrectly predicting its final position. Notably, o3-mini successfully predicts the correct output for the episodes in Figures 6 and 7, but fails on the example in Figure 5. Figures 8 and 9 highlight episodes from the systematicity task. As shown, all general-purpose LLMs fail to produce accurate transformations, often misplacing the transformed object within the grid. In contrast, the model trained via MLC consistently predicts the correct transformation.
@@ Appendix E Additional results
As outlined in Section D.2, the general-purpose LLMs we evaluate are instructed to present their final output grid predictions using the keyword “output:”, followed by a two-dimensional array of size $10\times 10$ in the format $\ldots$]]”. Responses that do not conform to this expected pattern are excluded from subsequent analyses and are not included in accuracy calculations. Table 4 provides an overview of the proportion of valid responses for each model. In the standard few-shot learning setting, all models demonstrate very high valid response rates, exceeding 99%. However, in the systematicity task, a slight decrease in valid responses is observed for Gemini 2.0 Flash when additional visual input (text+image) is introduced, with the rate falling to 94.09%. More significantly, GPT-4o exhibits a notable drop in valid response rate to 77.24% under multimodal conditions. We hypothesize that this decline may be attributed to the increased context length resulting from the additional image input.
Error Analysis.
As described in Section 5.2, we analyze the models’ predictions and compare them with common failure modes. Table 5 shows the percentage of each error type described in Section 5.2 across models. For errors related to the models predicting a primitive or level-1 transformation instead of the desired level-2 transformation composition, we further illustrate which specific primitive or level-1 transformation was applied in Table 6. Specifically, this table shows whether the primitive transformation applied was based on the object’s shape, color, or neighboring object. Similarly, the table illustrates which specific level-1 transformation composition was applied.
Table 5: Error distribution by error category across models. Values denote the percentage (%) of prediction errors assigned to each error category.
Model Format No Transform Primitive Level-1 Invalid Position Invalid Shape Other
GPT-4o 0.60 0.46 4.59 7.71 6.62 79.26 0.77
   + image 22.91 3.10 4.19 4.09 4.26 59.84 1.60
Gemini 2.0 Flash 0.26 1.56 11.41 22.41 5.73 58.32 0.30
   + image 7.60 0.72 9.05 15.52 4.60 61.83 0.68
o3-mini (low) 0.00 5.06 30.86 13.08 0.79 49.31 0.91
Llama-3.2-3B-ReARC 0.00 55.28 25.13 4.77 0.30 14.47 0.06
   + test-time training 0.00 0.44 6.18 67.13 10.72 15.52 0.00
Mistral-NeMO-Minitron-8B-Full 0.00 0.54 53.41 28.89 0.12 17.03 0.01
   + test-time training 0.00 2.47 5.05 71.55 13.31 7.62 0.00
MLC (ours) 0.05 3.34 3.56 9.11 3.02 70.57 10.35
@@
In addition to the model trained via MLC on a stream of dynamically changing visual interpretation grammars, as described in Section 3.2, we adopt the approach of Lake (2019) and train a transformer-based encoder-decoder on a dataset governed by a fixed visual grammar (referred to as basic seq2seq). This means that the indicator-transformation mappings are static across the whole dataset. For instance, if yellow objects translate one step downward, then this applies to all data samples across the dataset. Instead of episodes with few-shot examples, this dataset comprises individual input-output grid pairs, where the objective is to predict the output grid corresponding to a given input grid. This more closely resembles a standard training approach.
Table 6: Percentages of errors falling into each primitive and level-1 transformation error category.
Primitive Transformations Level-1 Transformations
Model Shape Color Neighbor Shape+Color Shape+Neighbor Color+Neighbor
GPT-4o 2.09 1.56 0.93 4.68 2.10 0.92
   + image 1.97 1.57 0.66 2.65 1.01 0.42
Gemini 2.0 Flash 4.22 4.70 2.49 16.17 4.05 2.19
   + image 2.96 3.66 2.43 9.58 3.45 2.49
o3-mini (low) 16.67 13.07 1.12 11.16 0.99 0.93
Llama-3.2-3B-ReARC 14.22 6.48 4.43 2.80 1.74 0.24
   + test-time training 0.27 5.16 0.76 8.81 35.68 22.64
Mistral-NeMO-Minitron-8B-Full 39.44 7.46 6.50 15.30 13.49 0.11
   + test-time training 0.05 4.08 0.91 8.05 38.81 24.69
MLC (ours) 0.16 1.08 2.32 5.98 2.16 0.97
@@
The model is trained for 200 epochs on the dataset using the parameters specified in Section C. While it successfully fits the training data (with an accuracy of over 99%), it fails to generalize to the out-of-distribution test set, achieving a test accuracy of 0.0%. This demonstrates that traditional model training, sample by sample, does not encourage systematic generalization to unseen compositions. Instead, systematicity requires a training procedure with examples over dynamically varying interpretation grammars, as described in Section 3.2.
Appendix F Use of AI assistants
@@
Table 8: Statistics of the dataset version including more diverse transformations. Listed are the number of episodes in the training, validation, and test sets. Additionally, the final query transformation compositions (level 2) are reported for both the training and evaluation datasets. The rightmost column details the frequency of each basic geometric transformation present in the training dataset.
Set No. Type Composition Transformation Freq.
seed 1860 Train 85528 Train translation+reflection+coloring red coloring 18376
Val 5472 reflection+rotation+extension orange coloring 18627
Test 5473 translation+reflection+rotation yellow coloring 18961
translation+rotation+coloring green coloring 18491
reflection+coloring+extension 1-step left translation 6471
reflection+rotation+coloring 2-step left translation 3671
translation+coloring+extension 1-step right translation 7942
rotation+coloring+extension 2-step right translation 5438
Test translation+rotation+extension 1-step up translation 6780
translation+reflection+extension 2-step up translation 4051
1-step down translation 6686
2-step down translation 4022
leftward extension 11742
rightward extension 12071
upward extension 11801
downward extension 12909
cw. rotation 20797
ccw. rotation 20799
horizontal reflection 23536
vertical reflection 23413